How would I go about solving this question?
A construction company pays $12,000 for a truck. The truck depreciates at a constant rate of 20% yearly. How long will it take for the truck to be worth $7200.
Thanks in advance.

Answer by Boreal(15235)

7200=12000(1-.0.20)^n, where n is the number of years.
0.6=0.80^n
logs of both sides
-0.2218=n log 0.80=n*-0.0969
divide both sides by -0.0969
=2.29 years
Rough check
after 1 year, truck is worth $12000*0.8=$9600
after 2 years, it is worth $9600*0.8=$7680
after 3 years, it is worth $7680*0.8=$6144. The answer above is reasonable.